### v1

# How Does a Deep Neural Network Look at Lexical Stress?

Abstract
Despite their success in speech processing, neural networks often operate as black boxes, prompting the question: what informs their decisions, and how can we interpret them? This work examines this issue in the context of lexical stress. A dataset of English disyllabic words was automatically constructed from read and spontaneous speech. Several Convolutional Neural Network (CNN) architectures were trained to predict stress position from a spectrographic representation of disyllabic words lacking minimal stress pairs  (e.g., initial stress WAllet, final stress exTEND), achieving up to 92% accuracy on held-out test data. Layerwise Relevance Propagation (LRP), a technique for CNN interpretability analysis, revealed that predictions for held-out minimal pairs (PROtest vs. proTEST) were most strongly influenced by information in stressed versus unstressed syllables, particularly the spectral properties of stressed vowels. However, the classifiers also attended to information throughout the word. A feature-specific relevance analysis is proposed, and its results suggest that our best-performing classifier is strongly influenced by the stressed vowel’s first and second formants, with some evidence that its pitch and third formant also contribute. These results reveal deep learning’s ability to acquire distributed cues to stress from naturally occurring data, extending traditional phonetic work based around highly controlled stimuli.
I Introduction
While deep learning based architectures have transformed many aspects of speech processing, those models often exhibit an intransparent prediction process, obscuring how specific decisions are made. This work explores the application of one technique for shedding light on such architectures through the lens of a well-studied phenomenon in speech science and speech processing: lexical stress. Lexical stress refers to the prominence, accentuation, or emphasis of a particular syllable within a word. In many languages, lexical stress is crucial for communication as it can can serve as the primary basis for a contrast in meaning. For example, in English, the words “PROtest” (noun; stress on initial syllable) and “proTEST” (verb; stress on final) have contrasting stress patterns but the same phonemes (with predictable changes to the quality of the initial vowel depending on stress).
A long tradition of phonetics research has examined the (relative) temporal, spectral, and energetic properties of stressed and unstressed syllables [5, 6]; see Cutler [4] and Gay [7] for reviews. Building on this work, automatic stress detection methods have used acoustic features such as fundamental frequency, amplitude, and energy [19, 29]. Bayesian classifiers with multivariate Gaussian distributions have been applied to enhance stress detection accuracy, utilizing acoustic features such as peak-to-peak amplitude, normalized energy, and vowel duration [30, 31]. Similar to other areas of speech processing, recent work using deep learning architectures has significantly improved over these more traditional approaches. Convolutional Neural Networks (CNNs) and Multi-Distribution Deep Neural Networks (MD-DNNs) outperformed traditional models in both English and Arabic stress detection in second language speech [25, 18]. Additionally, an attention-based neural network combined with data augmentation through a neural network based text-to-speech system has been introduced to detect lexical stress errors in second-language English learners, achieving significant improvements in precision and recall [15].
Although these advancements in deep learning have led to significant improvements in speech processing, they also introduce challenges, particularly in understanding classifier decisions and gaining insight into how these models operate. This issue of interpretability is crucial. Better understanding the acoustic features that drive model predictions can help us verify that the model will be reliable when applied to novel data (which may have different distributions for critical features). Neural network interpretability has broadly become an active research area, with a number of techniques being explored across different domains. Here, we focus on Layerwise Relevance Propagation (LRP) [2], a technique that was developed to provide insights into the decision-making processes of these complex models by understanding what specific portions of the input influence the model’s output. We use this technique to investigate how CNNs – which have demonstrated superior performance in lexical stress detection – derive their predictions from the spectral features in their input.
This paper offers several novel contributions. First, a new speech dataset was automatically gathered without human annotation, using a pipeline built around deep learning systems. Second, we present a state-of-the-art lexical stress classifier and show its ability to generalize to novel words, including words from previously unseen languages. Lastly, we employ LRP to analyze the spectral features these classifiers use to achieve high accuracy.
This paper is organized as follows. In the next section, we describe how the dataset was generated. In Section III we present how we built the classifier and how it was trained. In Section IV we present the LRP technique, its application in our study, and our techniques for analyzing LRP results. In Section V we empirically evaluate the classifiers. The LRP analyses are presented in Section VI, and the paper concludes with a discussion in Section VII.
Figure 1: End-to-end workflow for dataset construction, model training, and interpretability analysis. The blue-colored section illustrates the dataset creation and training process. Audio is collected from various datasets and extracted based on word- and phoneme-level timestamps. Words are tagged for stress using either part-of-speech tagging (minimal pairs) or dictionary entries (no minimal pair words). Once the audio is extracted, the training data (no minimal pair words) is expanded through data augmentation techniques to make CNN-based model training more robust. The red-colored section illustrates the processes of model interpretability and acoustic feature relevance measurement, applied to the test set of minimal pairs. Audio is converted into a spectrogram. This is processed by the trained model; LRP methods are applied to generate heatmaps, revealing the regions of the spectrograms that contributed most to the model’s predictions. The contribution of different (sub)syllabic regions to the LRP heatmap is quantified with IOU metric. For the stressed vowel, acoustic features are extracted and used to generate feature-specific heatmaps; these are compared to the full heatmap to determine which acoustic features are most dominant in the model’s decision-making process.
II Dataset Construction
Previous studies on lexical stress often depended on stress labels assigned by human annotators [25, 15, 29] or were assigned based on the canonical stress for the orthographic word taken from pronunciation dictionaries such as the CMU dictionary [18]. However, manual annotation is costly and can introduce inter-annotator inconsistency, and the use of a single citation form can yield errors for orthographically ambiguous minimal pair words (e.g., protest). Prior work also failed to explicitly document measures to prevent lexical overlap between training and evaluation splits, leaving open the possibility that identical word types were seen during both stages. Reported accuracy may therefore overestimate true generalization. To avoid these concerns, we created a novel dataset, using a fully automatic pipeline, without utilizing human annotations.
II.1 Selection of English Disyllabic Words
We prompted ChatGPT 4o [10] to generate 30 pairs of disyllabic words that form stress minimal pairs: pairs of words with the same spelling but different meaning depending on the primary stress placement. Four of these items could not be incorporated into other parts of our pipeline (the part-of-speech tagger, discussed below, failed to correctly annotate these forms). The remaining 26 pairs are address, conduct, content, contrast, convert, decrease, digest, entrance, export, extract, import, increase, insult, object, perfect, present, progress, project, protest, rebel, record, refuse, reject, subject, suspect, and transfer; the initial and final syllable stress variants of these words will be referred to as “minimal pairs.” We also asked ChatGPT 4o to generate a larger set of 250 disyllabic words that do not have stress minimal pairs (e.g., initial stress WALLet, final stress exTEND). After excluding one error, we retained 249 words (124 initial stress, 125 final stress). We refer to this as the set of “no minimal pair words.”
II.2 Extraction of Words from Existing Corpora
These words were then extracted from three existing English speech corpora. The first dataset is LibriSpeech [22], a widely-used corpus consisting of approximately 1,000 hours of read speech derived from audiobooks. The second dataset is The Supreme Court corpus [28], which includes recordings of oral arguments made before the United States Supreme Court. This dataset provides speech data from a highly formal and legal context, offering variation in speaker formality and legal terminology that can influence acoustics associated with lexical stress. The third dataset is TED-LIUM [9], derived from TED Talks. It includes over 400 hours of speech by speakers from various language backgrounds and nationalities, discussing a wide range of topics. The diversity in speaking styles and contexts in TED-LIUM further enriches the phonetic variation in our dataset.
The processing pipeline is shown in Figure 1. First, we used the Montreal Forced Aligner [21] to align the transcript with the audio, extracting precise time intervals for each target word in our minimal pairs and no minimal pair words lists in each of the datasets. After identifying the location of target minimal pair words within the audio, we applied the SpaCy [1] part-of-speech tagging system to the entire audio in order to determine whether each word was used as noun or verb. This classification follows linguistic conventions for English disyllabic words, where nouns typically carry primary stress on the initial syllable, referred to as initial stress (IS), and verbs carry primary stress on the final syllable, referred to as final stress (FS) [19]. For the no minimal pair, we used annotations from the CMU Pronouncing Dictionary [18], as the stress placement is unambiguous in this case.
Words were extracted precisely at the forced aligner start time and lasted 0.5 seconds. This fixed width was necessary for the use of the CNN architecture; we verified that this window was sufficiently large to encompass both syllables of every disyllabic word. We also split each minimal pair word into syllables using an automatically syllabified version of the CMU Pronouncing dictionary  [^1] The syllabified dictionary and syllabifier can be accessed at http://sourceforge.net/p/p2tk/. and phoneme-level time stamps from the forced aligner.
Overall, we had 6,659 samples (124 types) and 3,095 samples (125 types) for initial and final stress in the no minimal pairs set, respectively, and 5,547 samples (40 types) and 1,260 samples (40 types) for initial and final stress in minimal pairs, respectively.
II.3 Phonetic Properties of Extracted Samples
To provide an initial verification that our procedures yielded sets of disyllabic words that contrasted in stress location, two well-known correlates of lexical stress were examined: amplitude (vowels in stressed syllables are louder than those in unstressed syllables) and duration (vowels in stressed stressed syllables are longer than those in unstressed syllables) [5, 6]. For each sample, we calculated the ratio of the properties of the vowels of the initial versus final syllables. These relative measures allow us to control for potential differences in speech rate and amplitude across samples. Our samples reliably showed the expected differences. Samples with initial stress had high ratios of amplitude (mean 0.6, s.d. 0.18) and duration (mean 0.53, s.d. 0.15), reflecting relatively loud, long vowels in stressed initial versus unstressed final syllables. In contrast, samples with final stress showed lower mean ratios for amplitude (mean 0.29, s.d. 0.14) and duration (mean 0.35, s.d. 0.12), reflecting relatively quiet, short vowels in unstressed initial versus stressed final syllables. To confirm that this is a reliable difference between sample types, we used bootstrap resampling (with 1,000 replicates) to estimate the 95% confidence interval for differences in mean ratios between initial versus final stress samples. These differences were significant (amplitude mean difference: 0.305; 95% CI (0.302,0.309); p<p<.0001; duration mean difference: 0.178, 95% CI (0.175,0.18), $p<$ .0001). This provides independent confirmation that the samples in this data set have contrasting stress patterns.
III Lexical Stress Detection
III.1 Model Architectures
In our study, we investigated different CNN architectures to develop a binary classifier for identifying the initial or final stress of stress-minimal word pairs. The input to all models consists of magnitude spectrograms obtained using the Short-Time Fourier Transform (STFT), following z-score normalization. The STFT was computed with a sampling rate of 16,000 Hz, using a Hamming window of size 0.02 seconds and a window stride of 0.01 seconds. We experimented with several well-known CNN architectures. The LeNet-5 model consists of 3 connected convolutional layers with average pooling between the layers, followed by 2 dense layers in a 2x2 configuration [17]. The VGG11, VGG16, and VGG19 architectures consist of 8, 13, and 16 convolutional layers, respectively, with max pooling between the layers, followed by an additional 3 dense layers [27]. Finally, the ResNet18 model consists of 18 convolutional layers with residual connections to address the vanishing gradient problem during training [8].
Figure 2: Spectrograms and corresponding attributions for the words “PREsent” and ‘preSENT” using the VGG16 model. Orange vertical lines represent phoneme boundaries (the label on each line denotes the segment to the left of the boundary; 1 and 0 denote stressed and unstressed vowels, respectively). The blue vertical line represents the end of the initial syllable. Panel (a) shows the spectrogram of “PREsent” (IS), while panel (e) shows the spectrogram of ”preSENT” (FS). Panels (b)-(d) and (f)-(h) display attributions using various methods: $\mathbf{LRP_{\epsilon}},\mathbf{LRP_{\alpha 1}}\text{ and }\mathbf{LRP_{CMP}}$ respectively, shown over the spectrograms. These LRP variants effectively capture initial versus final stress contrasts, highlighting features at different scales. The $\mathbf{LRP_{\alpha 1}}$ (panel (c) vs. (g))and $\mathbf{LRP_{CMP}}$ methods (panel (d) vs. (h)) show clearly different patterns depending on the location of stress.
III.2 Data Splitting and Training Process
The data were partitioned so that no word type appears in more than one split, thereby eliminating lexical overlap between training, validation, and test sets.
The training set was constructed using 201 out of the 249 word types from the no minimal pair words. Data augmentation techniques were employed to enlarge this set of examples. For each sample, we added a low-pass filtered version of the sample (cutoff frequency at 3000 Hz) as well as three versions of the sample mixed with noise. A sample of multi-talker babble from the VOiCES Corpus [23]) was combined with the original sample at three different levels (20 dB, 10 dB, and 3 dB SNR). These augmentations aimed to simulate real-world conditions, enhancing the robustness of our models. After the augmentation, we had 29,810 and 13,510 samples from the initial and final stress, respectively, corresponding to the 201 no minimal pair word types. The augmented dataset will be referred to as the “no minimal pairs Train.”
The remaining 48 no minimal pair word types were divided into two lists of 24 words each, with one list used for the validation (i.e., hyper-parameter tuning;“no minimal pairs Validation”) and the other for the test set (“no minimal pairs Test”). The 26 minimal pairs words will be referred to as “minimal pairs Test”. Note that two test sets were used (no minimal pairs Test and minimal pairs Test) as the minimal pairs set may contain errors; stress was determined solely by part-of-speech tagging, which may be inaccurate. In contrast, the tagging for the no minimal pair words is more reliable; as the word form is unambiguous, the lexically-determined stress pattern can be accurately identified.
The models were trained using standard procedures for CNNs. To stabilize training, we employed an Adam Optimizer [12] with a learning rate schedule and batch normalization. To address class imbalance (i.e., more initial versus final stress examples), we employed focal loss [20], modifying the standard cross-entropy loss to focus more on hard-to-classify examples (the gamma parameter was tuned to 2 based on validation set performance).
Figure 3: Further illustrative examples of spectrograms and corresponding attributions for the words “subject,” “record” and “refuse” with different primary stress location, using the VGG16 model. Vertical lines and annotation follow Figure 2. Panel (a)-(d) show spectrograms and corresponding $\mathbf{LRP_{CMP}}$ heatmaps for the word ”subject” with initial and final stress, respectively. Similarly, panels (e)-(l) display spectrograms and heatmaps of the words “record” and “refuse,” with different stress types.
IV Interpretability Analysis Methods
IV.1 Layer-wise Relevance Propagation
Layer-wise Relevance Propagation (LRP) has been extensively used in computer vision to provide insights into the contribution of individual input features to model predictions [13]. Starting at the output, LRP attributes the classification prediction to units in the preceding layer – how relevant is each unit to the classifier producing one output versus another? This process is then repeated, back-propagating the relevance score of each unit to the classification prediction until the input layer is reached. This technique allows us to attribute relevance scores to each feature (e.g., how much does pixel X contribute to classifying an image as dog versus cat?).
In our study, we apply this technique (as implemented in the Captum library [14]) to our lexical stress classifiers, backpropagating prediction to time-frequency bins in the input. This method identifies the regions of the spectrogram that have the most significant influence on the model’s classification of lexical stress. We specifically examine our best-performing VGG-based classifier, VGG16, and ResNet18 (for model performance comparisons see Section V and Table 1).
Denote by $\mathbf{x}^{l}$ the input to the $l$-th layer. Denote by $\mathbf{W}^{l}$ the weights associated with layer $l$. Denote by $\mathbf{Z}^{l}$ is the elementwise preactivation of layer $l$ and it is defined as $Z_{ij}={W}^{l}_{ij}\odot{x}^{l}_{i}$. The relevance score of layer $l$ is computed recursively from a higher layer $l+1$. Relevance score $\mathbf{LRP_{Z}}$ is defined as [13]
$R_{i}^{l}=\sum_{j}\frac{Z^{l}_{ij}}{Z^{l}_{j}}R_{j}^{l+1}~,$ (1)
where $Z^{l}_{j}=\sum_{i}Z^{l}_{ij}$. The $\mathbf{LRP}_{Z}$ method effectively traces the influence of each element of the input through the network, providing insight into how these impact the model’s predictions.
In addition to this method, we experimented with several types of LRP to determine the most effective method for identifying the regions of the spectrogram that are most influential for classifying words as stressed versus unstressed. The $\mathbf{LRP_{\epsilon}}$ rule extends the basic LRP method by addressing challenges such as division by zero and the influence of weak or noisy preactivations. It is defined as:
$R_{i}^{l}=\sum_{j}\frac{Z^{l}_{ij}}{Z^{l}_{j}+\epsilon\,\mathrm{sign}(Z^{l}_{j})}R_{j}^{l+1}~,$ (2)
In this rule, a small constant $\epsilon$ is added to the denominator along with the sign of $Z^{l}_{j}$. This modification prevents division by zero and mitigates the impact of small or noisy preactivations $Z^{l}_{ij}$ on the relevance distribution. By incorporating $\epsilon$, the $\mathbf{LRP}_{\epsilon}$ rule enhances the stability of relevance propagation, leading to more reliable and interpretable results. This adjustment is particularly beneficial when dealing with minimal or problematic contributions, as outlined by Bach et al. [2].
The $\mathbf{LRP_{\alpha\beta}}$ rule, introduced by Bach et al. [2], is an extension of the basic LRP method that differentiates the contributions of facilitatory and inhibitory activation flow. It is defined as:
$R_{i}^{l}=\sum_{j}\left(\frac{\alpha Z_{ij}^{+}}{Z_{j}^{+}}+\frac{\beta Z_{ij}^{-}}{Z_{j}^{-}}\right)R_{j}^{l+1}$ (3)
where $Z_{ij}^{+}=\max\{Z^{l}_{ij},0\}$ and $Z_{ij}^{-}=\min\{Z^{l}_{ij},0\}$. In this rule, $\alpha$ and $\beta$ are non-negative parameters that weight the relevance distribution for facilitatory and inhibitory preactivations, respectively. The sum of $\alpha$ and $\beta$ is constrained to 1 to ensure the conservation of relevance between layers. This approach allows for a more nuanced allocation of relevance based on the nature of the preactivations.
Earlier implementations of LRP, such as those described in Bach et al. [2] and Lapuschkin et al. [16], typically employed a single LRP rule applied uniformly throughout the network. This approach failed to account for the varying needs of different layers or components within the network, resulting in less accurate and insightful relevance attributions. To address these limitations, Kohlbrenner et al. [13] introduced a composite strategy, which employs a combination of multiple LRP rules tailored to different parts of the network. In our analysis, an identity function was used to backpropagate activation for the first two convolutional layers closest to the input, $\mathbf{LRP}_{\alpha 1}$ ($\mathbf{LRP}_{\alpha\beta}$ with $\alpha=1$) for the remaining convolutional layers, and $\mathbf{LRP}{\epsilon}$ for the fully connected layers closest to the output. This combined approach, denoted as $\mathbf{LRP_{CMP}}$, enhances model interpretability and provides more accurate relevance attributions.
[…]
Figure 4: Heatmaps of the first two formants ($F_{1}$ and $F_{2}$) and the original heatmap generated by $\mathbf{LRP_{CMP}}$ using VGG16 for the word ”PREsent”. Panels (a) and (b) display the heatmaps generated for the two formants, considering bandwidth, spectrogram intensity, and voice intensity within the relevant time points. Finally, panel (c) illustrates the heatmap attributes generated by $\mathbf{LRP_{CMP}}$ utilizing VGG16. Orange vertical lines represent phoneme boundaries, and the blue vertical line represents the end of the initial syllable, same as Fig. 2.
IV.2 Intersection over Union in Sub-Syllabic Regions of the LRP heatmap
We quantitatively assess the LRP analysis to assess how regions corresponding to different components of the initial and final syllables contribute to lexical stress classification using the Intersection Over Union (IOU) metric, adapted for analyzing relevance attribution in heatmaps. Specifically, we define the inside-total overlap ratio $\mu$:
$\mu=\frac{R_{in}}{R_{tot}}$ (4)
Here, $R_{in}$ is the sum of all pixels of the spectrogram inside a bounding box derived from LRP heatmap, and $R_{tot}$ is the total sum of relevance values across the entire heatmap. This metric $\mu$ effectively measures the proportion of relevance attributed to the area of interest, providing insight into what portions of the signal the model has paid attention to. Higher values of $\mu$ indicate that a larger fraction of the relevance is concentrated within the target area.
The bounding boxes defining the regions of analysis were created using the phoneme-level timestamps and syllable boundaries described above, ensuring alignment with the relevant parts of the spectrograms. We set bounding boxes around the initial and final syllables and then, within each syllable, separate the vowel region from any other portion of the syllable (onset/coda).
IV.3 Feature-Specific Relevance
In addition to the IOU metric, we introduce a measure to evaluate how the model focuses on specific acoustic features during decision-making. Focusing on the vowel within the stressed syllable, this analysis aims to isolate the contribution of each acoustic feature to the model’s decision-making process. We constructed feature-specific heatmaps that represent the relevance distribution as if the model were focusing exclusively on the given feature. This was done by extracting the fundamental frequency ($F_{0}$), the first three formants ($F_{1},F_{2}$ and $F_{3}$) along with their formant bandwidth and sound intensity using the default settings of the Praat analysis software [3]. For each formant, heatmap values at each time point were generated by normalizing the intensity within the formant’s bandwidth and then scaling these normalized values by the sound intensity.
To quantify the similarity between these feature-specific heatmaps and the original relevance heatmap produced by LRP, we employed the Pearson correlation coefficient $r$. By comparing the strength of the correlation between feature-specific heatmaps, we gain insight into the relative importance of each acoustic feature in the model’s decision-making process. Additionally, we examine the ability of heatmaps created by all possible combinations of these features to determine how these different acoustic features interact to contribute to the model’s predictions.
Finally, we also examine the limitations of these specific features by considering the distribution of residual pixels left unexplained by any of these features ($F_{0}$ to $F_{3}$). We mask the LRP heatmap using the heatmap generated by combining all these features. We then examine the relative distribution of these unexplained pixels across the spectrum of the stressed vowels to consider what additional features, beyond pitch and formants, contribute to model performance.
V Model Performance
Classifier results on the test sets are summarized in Table 1. The VGG16 architecture achieved classification accuracies of 87% and 92% in distinguishing between initial and final stress in minimal pair and no minimal pair words test sets, respectively. The lower accuracy in the minimal pair test set may reflect errors in part-of-speech tagging. Note as well that human perception of lexical stress in disyllabic words based solely on acoustics (without access to syntactic or other contextual cues) is not at ceiling; [32] reported first-language English listeners had a mean accuracy of 84% in this task.
Other architectures also showed strong results. VGG19 matched VGG16’s performance, despite utilizing a more complex architecture, suggesting that VGG16 is sufficient. VGG11 showed slightly lower performance on both test sets. While the other architectures also showed slightly lower performance, they still achieved a high level of performance (>> 85% on both test sets).
Table 1: Classification accuracy for each model architecture on each test set.
Test Set Classifier Accuracy (%)
Minimal pair LeNet 85
VGG11 86
VGG16 87
VGG19 87
ResNet18 86
No minimal pair LeNet 90
VGG11 91
VGG16 92
VGG19 92
ResNet18 90
We also examined the ability of the classifier to generalize to other languages. While there is considerable cross-linguistic variation, there is also some degree of similarity in the acoustic realization of lexical stress across languages [4], leading us to expect some degree of generalization. We therefore examined the performance of the trained VGG16 and ResNet18 models before and after fine-tuning on both Dutch [24] and Hebrew [26] datasets. VGG16 achieved 93% accuracy on the Dutch data and 88% on the Hebrew data after fine-tuning. Similarly, ResNet18 achieved 93% on Dutch and an accuracy of 85% on Hebrew. This suggests our modeling framework provides a robust approach to detecting lexical stress. The results are summarized in Table 2.
Table 2: Classification accuracy for other languages.
Test Set Classifier Accuracy (%)
Dutch VGG16 81
ResNet18 79
VGG16 after fine-tuning 93
ResNet18 after fine-tuning 93
Hebrew VGG16 68
ResNet18 73
VGG16 after fine-tuning 88
ResNet18 after fine-tuning 85
VI Interpretability Results
We utilized LRP and our interpretability metrics to examine more closely the aspects of the input that influenced the strong performance of these classifiers.
VI.1 Layer-wise Relevance Propagation (LRP) Analysis
Inspection of the LRP heatmaps for both initial and final stress words suggested they highlighted distinct areas of relevance that align with the energy distribution patterns observed in the spectrograms. Figures 2 and 3 illustrate this for the best-performing VGG16 classifier. For IS words, the LRP heatmaps emphasized the initial syllable as the most relevant area for classification by VGG16. Conversely, for FS words, the LRP heatmaps showed a concentration of relevance towards the end of the phoneme sequence. Figure 3 provides further illustrations of these differences, focusing on the $\mathbf{LRP_{CMP}}$ method.
VI.2 Bounding Box and IOU Analysis
The IOU measurements provide a quantitative assessment of the overlap between the bounding boxes and the relevant pixels in the LRP heatmaps.
Table 3 shows IOU measurements ($\mu$) for different LRP methods, examining VGG16 and ResNet18. The $\mathbf{LRP_{CMP}}$, $\mathbf{LRP_{\alpha 1}}$, $\mathbf{LRP_{\epsilon}}$ methods suggest that both classifiers attend primarily to acoustic information in stressed versus unstressed syllables, and that the greatest concentration of pixels is specifically in the stressed vowel region. The basic $\mathbf{LRP_{z}}$ method results are mostly consistent with these generalizations, but give some greater weight to unstressed syllables.
Overall, these findings suggest that the classifiers predict the location of stress by attending to features of the stressed syllable, in particular to stressed vowels. However, it is important to note that under all methods, a considerable number of pixels lie outside of the stressed vowel – consistent with the distribution of cues to lexical stress throughout the entire syllable. Furthermore, many pixels lie outside of the stressed syllable – consistent with a view that sees lexical stress as a “relational” property of stressed as compared to unstressed syllables.
Table 3: Average IOU values for different classifiers and LRP methods, separated by stressed versus unstressed syllables and syllable region (vowel versus all other components of syllable).
Classifier LRP Method $\mu_{\mathrm{vowel}}$ $\mu_{\mathrm{not\,vowel}}$
Stress No stress Stress No stress
VGG16 $\mathbf{LRP_{CMP}}$ 0.432 0.135 0.276 0.156
$\mathbf{LRP_{\alpha 1}}$ 0.419 0.129 0.279 0.173
$\mathbf{LRP_{\epsilon}}$ 0.428 0.137 0.27 0.164
$\mathbf{LRP_{z}}$ 0.301 0.154 0.318 0.226
ResNet18 $\mathbf{LRP_{CMP}}$ 0.37 0.136 0.298 0.195
$\mathbf{LRP_{\alpha 1}}$ 0.361 0.128 0.255 0.255
$\mathbf{LRP_{\epsilon}}$ 0.34 0.13 0.275 0.254
$\mathbf{LRP_{z}}$ 0.326 0.12 0.204 0.364
Table 4: $r$ and $p$ values for heatmaps based on different features in the stressed vowel region (averaged across samples), using VGG16 and applying $\mathbf{LRP_{CMP}}$ method (sorted by $r$; higher $r$ and lower $p$ are better). See text for details.
Method Acoustic Feature Mean $r$ Mean $p$
VGG16  +  $\mathbf{LRP_{CMP}}$ $F_{1}$} 0.366 0.003
$F_{1},F_{2}$} 0.359 0.003
$F_{1},F_{2},F_{3}$} 0.305 0.002
$F_{1},F_{3}$} 0.277 0.001
$F_{0},F_{1},F_{2}$} 0.245 0.001
$F_{0},F_{1},F_{2},F_{3}$} 0.237 0.001
$F_{0},F_{1},F_{3}$} 0.205 0.001
$F_{0},F_{2}$} 0.154 0.003
$F_{2}$} 0.154 0.008
$F_{2},F_{3}$} 0.136 0.008
$F_{0}$} 0.116 0.004
$F_{0},F_{3}$} 0.114 0.003
$F_{3}$} 0.048 0.02
VI.3 Feature-Specific Relevance Analysis
We extended our IOU analysis by investigating how VGG16’s predictions are influenced by the phonetic features of the signal region that most strongly drives its responses, namely, the stressed vowels. We compared each of these features to the $\mathbf{LRP_{CMP}}$-derived heatmap, focusing on the heatmap region corresponding to the stressed vowel. Figure 4 shows two feature-specific heatmaps corresponding to $F_{1}$ and $F_{2}$ in isolation. Additionally, the original relevance heatmap produced by VGG16 and $\mathbf{LRP_{CMP}}$ is presented for comparison.
As shown in Table 4, a heatmap generated solely using $F_{1}$ of the stressed vowel exhibits the strongest correlation with the observed heatmap. Unsurprisingly, this formant makes a significant contribution, as it correlates with tongue height, jaw height, and the overall opening of the oral cavity, which in turn are strongly associated with vowel sound intensity. However, the combination of $F_{1}$ and $F_{2}$ is a close second to $F_{1}$ alone, suggesting that the two acoustic features traditionally related to vowel quality in English work together to determine classifier predictions. Similarly, heatmaps combining these two formants with $F_{3}$ and, to a lesser extent, $F_{0}$, are used to predict LRP responses. However, the outsize influence of $F_{1}$ is made clear by the low performance of the other formants and pitch in isolation.
While the $r$ values are significant, they suggest that these features, even when considered in combination, fail to explain many aspects of the LRP heatmap for stressed vowels. To explore this, we examine the frequency distribution of residual (unexplained) pixels. As shown in Table 5, the vast majority of unaccounted-for pixels lie in between $F_{0}$ and $F_{1}$.
What other acoustic features might explain these pixels? One possibility is that the classifiers are using this information to track pitch, a key contributor to lexical stress perception [4]. Our classifier relies on a standard spectrographic representation as input. The relatively poor resolution of this representation at low frequencies makes it a poor estimator of pitch. Tracking the spacing of harmonics (which have high amplitude in this region of the signal) would give the classifier access to this critical information. In fact, inspection of Figures 2 and 3 suggests that the classifier is attending to regularly spaced concentrations of energy in these lower frequencies – which could correspond to harmonics. This could allow the classifier to estimate pitch with a higher degree of resolution than is available in the region of the spectrogram corresponding to $F_{0}$, using information that can be readily uncovered from the spectrographic input.
Table 5: Relative distribution of pixels (averaged across samples) in the stressed vowel region of the residual $\mathbf{LRP_{CMP}}$ heatmap for VGG16 (see text for details).
Region Initial stressed Vowel Final stressed Vowel
Between $F_{0},F_{1}$} 0.93 0.886
Between $F_{1},F_{2}$} 0.07 0.1134
Between $F_{2},F_{3}$} 0 0
Above $F_{3}$} 0 0
VII Discussion
To explore what allows CNNs to outperform more traditional approaches, we built several novel classifiers for lexical stress detection, using LRP-based analyses to explore the spatiotemporal properties that drive model performance. A fully automatic pipeline was used to extract disyllabic words from read and spontaneous speech. CNNs trained on words without minimal pairs exhibited high performance on held-out no minimal pair words and words with minimal pairs. Follow-up analyses showed that with fine-tuning, the classifiers exhibited high accuracy on two other languages (Dutch and Hebrew). LRP analyses on the spectrogram input space for the CNNs revealed that the models were sensitive to acoustic properties in both stressed and unstressed syllables, but showed greater reliance on stressed syllables (particularly the stressed vowel). Analysis of our highest-performing model showed that $F_{1}$ of the stressed vowel exerted a stronger influence on predictions than other formants, but $F_{2}$ and, to a lesser extent, $F_{3}$ also contributed. Furthermore, these features failed to capture a substantial amount of variance in the LRP heatmap, suggesting additional features of the stresseed vowel contribute to its predictions (e.g., spacing of harmonics, particularly below $F_{1}$, may be used to track pitch).
The classifiers’ greater reliance on the properties of stressed versus unstressed syllables and more specifically stressed vowels versus other syllable regions is similar to the focus of the phonetics literature on stressed syllables and stressed vowels [4, 7]. Critically, differing from some early phonetics studies, the classifier does not focus exclusively on these regions, but also attends to information in both syllables, including non-vowel portions [^2] It is possible that distributed nature of stress cues reflects errors in our forced aligner. However, given the substantial temporal range of LRP pixels (see, e.g., Figure 4(c)), it’s likely that the results are robust to small changes to segment boundaries.. The use of information in the unstressed syllable resonates with the now-widespread use of relativized or normalized cues to lexical stress [4]. The use of information outside of the vowel is also consistent with phonetics research that examines whole-syllable duration and intensity [4]. More broadly, the use of consonantal information is consistent with the view that stress is not a property of vowels but of whole syllables, organized within more complex prosodic structures such as feet; these structres condition not only the properties of vowels but also surrounding consonants in stressed and unstressed syllables (see, e.g., [11]).
Our feature-specific relevance analysis allowed us to connect the LRP results to formants of the stressed vowel. However, these analyses left a great deal of the LRP signal unexplained – not only in the stressed vowel, but also in other regions of the stressed syllable and the entirety of the unstressed syllable. Future work examining what acoustic features these pixels correspond to may provide new insights into the phonetics of English stress.
More broadly, this work suggests that deep learning may offer an important complement to traditional approaches to phonetics. Rather than starting with highly controlled materials like minimal pairs elicited under laboratory conditions, our classifier was trained on “messy” data – uncontrolled materials, varying along many dimensions beyond the contrast of interest, drawn from speech produced by a number of talkers in a variety of conditions. When exposed to this (naturally occurring) variability, the classifiers were able to extract regularities that fit with the previous phonetics literature. With increasingly powerful interpretability tools, deep learning may help phoneticians better understand the acoustic properties associated with linguistic contrasts.
Acknowledgments
Supported by NSF DRL Grant No. 2219843 and BSF Grant No. 2022618.
Author Declarations
Conflict of Interests
The authors have no conflict of interests to declare.
Data Availability
The data will be made available in a public repository prior to publication. Code and a demo can be accessed here: https://github.com/ItaiAllouche/minimalPairsLexicalStress.